(how to enter it in BA2 plus calculator) I have $72,000 in my bank account today. The bank manager told me that I will have $100,000 money in five years if I keep the money in this account and don’t touch it (meaning that I won’t withdraw from it or don’t add more into it). The account pays me daily interest. What is the APR for my account?
Answer #1

Given that

Present value = $72000

final value after 5 years = $100000

assume 365 days in a year

total days to maturity = 5*365 = 1825

using financial calculator with values

N = 1825

PV = -72000

FV = 100000

PMT = 0

compute for I/Y, we get, I/Y = 0.018%

this is daily effective rate,

So, APR compounded daily = 365*0.018 = 6.57%